Prev Next

SW&IT Meeting Minutes 2022-07-18

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • xxx
  • This week
    • xxx

YZ

  • Previous:
    • TMS test on Wed with Heejoo and Olga.
    • TMS software work.
    • Some support work (SX adsec -> rigid update with Xianyu).
    • Some MODS utilities work (and discussions with Olga).
    • LBC code study and planning.
  • Next:
    • Mainly focus on TMS and LBC work.
    • Some misc MODS and support work.

DE

  • Last week:
    • Fighting with AzCam VM
    • Web stuff with CV; I apparently am Web Updater Person now
    • WebIO unit chat; I'll go in this week and put them in DT's office
  • This week:
    • JW offboarding. He has a meeting with MS to go over his workflow.
    • Office trip for Synology firmware update, other tasks

GE

  • Last week:
    • Released a new test version of the LUCI Observer GUI which launches the RTDs as subprocesses rather than just creating windows, ensuring that they and their Aladin instances don't interfere with each other or the main Observer GUI.
    • Continued investigating double-buffering by instrumenting the GEIRS service to trace the origin of commands. This led to further investigation of the LUCI startup sequence.
  • This week:
    • I will be working half days (mornings 8-noon) every day through July 22nd, and taking July 21st off.
    • Preparation for the LUCI Database cleanup (IT 8454).
    • IT 8571, 8529 (Observer Scheduler GUI).

BM

  • Last week:
    • Fix for anemometermon problem (wrong exception getting caught, wasn't closing the connection). Sent out test plan to aoops.
    • soul-dxwfs processes suddenly dying (8666), saw masterdiagnostic/housekeeperWFS segfaults. turned on core dumps for future.
    • ATT
    • Shark-NIR SW meeting, nothing needed until Fall
    • RTDB disconnects
  • This week:
    • RTDB disconnects (7213, 8397, 8628)
    • ITs (reviewing logs): 8307 (offsetxy), 8553/8250/3957 (bcu timeouts + ocam not live)

SH

  • 507 problems with SHARK-NIR connections, rewiring some Steward offices, switch count for UITS, Louise PC
  • New core switches for mountain
  • Surplus and inventory
  • ITGC stuff
  • Dan R. shutter-door-closing HDI panel install questions (short)

next:

  • another Steward trip later in the week
  • data move from old to new Synology in Tucson
  • check Tucson KVM (not accessible on network)
Topic revision: r6 - 18 Jul 2022, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback